OpenStudy (mimi_x3):

The polynomial P(x) = x^3 - 2x^2 + kx + 24 has roots a , b and γ (i) It is known that two of the roots are equal in magnitude but opposite in sign. Find the third root and hence find the value of k Can you please help .

OpenStudy (anonymous):

:@

OpenStudy (mimi_x3):

:@ ?? why yu angry ? LOL

OpenStudy (anonymous):

3rd root is 2

OpenStudy (anonymous):

not angry

OpenStudy (anonymous):

just having a bad day

OpenStudy (mimi_x3):

how is 3rd root 2 ?

OpenStudy (mimi_x3):

lols , having a bad day ? why ?

OpenStudy (anonymous):

I passed the entry test to uni, but apparently, I need to get minimum of 2 A levels to get into it. A levels take 2 years to get. I have 4 months.

OpenStudy (mimi_x3):

ohh ok LOL

OpenStudy (anonymous):

I got so far (for roots a,-a and c) 3a +c = 2 and a^2 c = 24 U get that?

OpenStudy (mimi_x3):

yeah

OpenStudy (mimi_x3):

w8 , what's c ?

OpenStudy (anonymous):

I don't mean understand, I mean u calculated same?

OpenStudy (anonymous):

I said for roots, a,-a and c

OpenStudy (mimi_x3):

ohh lols , i dnt even know how to do it

OpenStudy (anonymous):

There is similar (not same) formula like Vieta for cubic as for quadratic....

OpenStudy (mimi_x3):

What's Vieta ?

OpenStudy (anonymous):

Formulae for sum and product of roots of quadratic.

OpenStudy (mimi_x3):

ohh that

OpenStudy (anonymous):

Let a be the first root >0. Then b=-a Now since a cubic function with 3 real roots then must be of the form: (x-a)(x-b)(x-v)= x^3 - 2x^2 + kx + 24 But b=-a so LHS is (x-a)(x+a)(x-v)=(x^2-a^2)(x-v) Multiplying the left hand side out we get: =x^3-v*x^2-a^2*x+v*a^2=P(x) From this we see that v=2 which implies 24=2*a^2 then a^2=12 but -a^2*x=kx thus k=-12

OpenStudy (mimi_x3):

thank you (: